Hilltop Manufacturing uses a predetermined manufacturing overhead rate based on direct labor hours to allocate manufacturing overhead to jobs. Selected data about the company's operations follows:

Actual manufacturing overhead cost $500,900
Estimated manufacturing overhead cost $551,000
Estimated direct labor cost $175,200
Estimated direct labor hours 50,500
Actual direct labor hours 60,800
Estimated machine hours 40,300
Actual machine hours 35,800

By how much was manufacturing overhead overallocated or underallocated for the year? (Round intermediary calculations to the nearest cent.)     

A.

$162,428 overallocated     

B.

$50,100 overallocated

C.

$50,100 underallocated     

D.

$162,428 underallocated
Answer #1

predetermined manufacturing overhead rate=Estimated manufacturing overhead cost/Estimated direct labor hours

=551,000/50,500

=$10.91 per direct labor hours

Hence applied overhead=predetermined manufacturing overhead rate*Actual direct labor hours

=10.91*60800

=$663328

Hence since applied overhead is greater than actual manufacturing overhead;

Overallocated overhead=663328-500900

=$162428